OpenStudy (anonymous):

Two alternate interior angles formed by parallel lines cut by a transversal have measures of (2x + 10)º and (3x-15)º. What is the measure of each? (Hint: Sketch the problem before you solve it.)

OpenStudy (lgbasallote):

 according to the rule of transversals, alternate interior angles are congruent so... 2x + 10 = 3x - 15 solve for x first and tell me what you get

OpenStudy (anonymous):

x = 5?

OpenStudy (lgbasallote):

hmm no. 2x + 10 = 3x - 15 subtract 2x from both sides 10 = x - 15 add 15 to both sides x = 25 got it?

OpenStudy (anonymous):

so how do i find my angles?

OpenStudy (lgbasallote):

2x + 10 <--substitute 25 to x => 2(25) + 10 what is the result?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

thank you so much!

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

welcome